# Load Forecasting for Households and Energy Communities: Are Deep Learning Models Worth the Effort?

### v4

@@ Abstract
Energy communities (ECs) play a key role in enabling local demand shifting and enhancing self-sufficiency, as energy systems transition toward decentralized structures with high shares of renewable generation. To optimally operate them, accurate short-term load forecasting is essential, particularly for implementing demand-side management strategies. With the recent rise of deep learning methods, data-driven forecasting has gained significant attention, however, it remains insufficiently explored in many practical contexts. Therefore, this study evaluates the effectiveness of state-of-the-art deep learning models—including LSTM, xLSTM, and Transformer architectures—compared to traditional benchmarks such as K-Nearest Neighbors (KNN) and persistence forecasting, across varying community size, historical data availability, and model complexity. Additionally, we assess the benefits of transfer learning using publicly available synthetic load profiles. On average, transfer learning improves the normalized mean absolute error by 1.97%pt when only two months of training data are available. Interestingly, for less than six months of training data, simple persistence models outperform deep learning architectures in forecast accuracy. The practical value of improved forecasting is demonstrated using a mixed-integer linear programming optimization for ECs with a shared battery energy storage system. The most accurate deep learning model achieves an average reduction in financial energy costs of 8.06%. Notably, a simple KNN approach achieves average savings of 8.01%, making it a competitive and robust alternative. All implementations are publicly available to facilitate reproducibility. These findings offer actionable insights for ECs, and they highlight when the additional complexity of deep learning is warranted by performance gains.
@@ 1 Introduction
Energy communities (ECs) are continuously gaining attention, as they accelerate the transition to renewable energy and enhance self-sufficiency by enabling local balancing of supply and demand. ECs are groups of individuals or organizations that produce, manage, and consume energy locally, through the use of public grid infrastructure, often with the goal of increasing sustainability and reducing energy costs [1]. For an efficient operation of these entities—and to improve both self-sufficiency and financial benefit [2]—model predictive control (MPC) methods have been employed with notable success [3, 4].
@@
In May 2024, Beck et al. [20] introduced the xLSTM model. The authors claimed that this model performs favorably when compared to state-of-the-art Transformers, both in performance and scaling. They demonstrated that xLSTM exhibits performance comparable to Transformers, with notable advantages such as linear rather than quadratic growth in memory consumption as the model input size increases. This improvement addresses scalability challenges, making xLSTM a more efficient option to handle larger inputs. A first work by Kraus et al. [21] demonstrates the effectiveness of xLSTM for long-term forecasting across various publicly available time series, with forecast horizons ranging from 96 to 720 hours. In contrast to prior work, the present study applies xLSTM in a short-term forecasting context, using a 24-hour horizon aligned with the requirements of our MPC application.
ECs often lack sufficient historical records [22], which can degrade model performance, when training any ML-based model. To address this challenge, transfer learning emerges as a promising solution by enabling knowledge reuse from related data. This approach is based on the idea that knowledge learned during one task can be reused to boost the performance on a related task [23]. In Moosbrugger et al. [24], the potential of leveraging transfer learning was demonstrated using open-access synthetic (also called standard) load profiles to improve load forecast error for renewable ECs with limited historical data. The study utilized a bidirectional LSTM model pre-trained on open-access synthetic profiles, achieving a 62% reduction in mean squared error (MSE) for forecasted load in an energy community of 74 households, while also significantly improving training stability.
In comparison to the aforementioned forecasting studies, this work is broader in scope. Load profiles, aggregation levels, training data durations, and model sizes are varied, providing a more comprehensive evaluation. This approach aligns with the recommendation of Pinheiro et al. [25], who noted that many existing studies demonstrate the superiority of a proposed technique on narrowly defined datasets, often hiding their weaknesses. By introducing diverse setups, this study seeks to reflect realistic variations, providing insights that are both more generalizable and practically applicable. In addition, a case study is presented that quantifies the financial implications of different forecast errors. In the case study, MPC is applied to a simulated EC with different forecasting algorithms.
In summary, the key contributions of this paper are as follows:
@@
    The performance of state-of-the-art deep learning models in short-term load forecasting is evaluated, with the recently proposed xLSTM model compared against established approaches such as LSTM and Transformers.
@@
    A comprehensive sensitivity analysis is carried out, where the aggregation level of ECs, the availability of training data, and the model complexity are varied.
@@
    The impact of transfer learning on the forecast accuracy is systematically assessed.
  4. 4.
    The financial implications of the load forecast accuracy regarding a flexibility optimization are illustrated within a case study.
Therefore, this study enhances current knowledge on the suitability of state-of-the-art deep learning algorithms for load forecasting in households and ECs, by assessing model sensitivity to common uncertainties and quantifying the impact and value of accurate forecasting in these settings.
The remainder of this study is organized as follows: In Section 2, we introduce the forecasting framework designed for ECs and give insight into the financial optimization case study. The results, discussed in Section 3, provide practical guidance for model selection based on forecast accuracy, model complexity, and economic impact. Section 4 concludes our study and outlines directions for future work.
@@ 2 Methods
In the next step, three deep learning models—xLSTM, LSTM, and Transformer—were compared against two baseline models: Persistence Prediction and k-Nearest Neighbors (KNN). Each model performs day-ahead short-term load forecasting, predicting an hourly power load profile from midnight to the following midnight. The models were trained and tested on various configurations in terms of community size, training dataset size, and model complexity. The practical value of improved forecasting was analyzed through a mixed-integer linear programming (MILP) optimization model for ECs with shared battery energy storage.
@@ 2.1 Data description
This study is based on 4,500 household load profiles from the UK Power Networks smart meter dataset [28], which contains electricity consumption data from households in London. To construct virtual ECs, subsets of 1, 2, 10, 50, or 100 households were randomly selected from the dataset. This selection process was repeated 20 times for each community size, without replacement, resulting in $100$ unique virtual ECs. All profiles span the period from July 2012 to February 2014, providing a robust foundation for the analysis. Q4 of 2013 was used as the test period, with the preceding 12 months used for training in all baseline model evaluations. The effect of using different quarters of 2013 as the testing period was also evaluated in a robustness analysis, which can be found in B.
@@ 2.2.1 Input Features
  • Day of the week: Encoded as a one-hot vector, represented as a binary 7-dimensional vector with one ”hot” bit, with public holidays assigned to Sunday [14, 31].
@@ 3.2 Sensitivity Analysis
It is seen that the nMAE of the persistent predictor is smaller compared to the deep learning models with up to six months of training data. However, with nine months of training data or more, the deep learning models outperform persistent prediction, with a difference in nMAE ranging from 1.8%pt to 2.0%pt, depending on the model type. This finding is particularly valuable for newly established ECs, which can initially rely on persistence prediction for MPC and transition to more sophisticated models after operating for nine months.
@@
As illustrated in Fig. 9(a), transfer learning using synthetic load profiles shows a notable reduction in nMAE for two months of available training data. Across all trained models, the average nMAE improvement is 1.97%pt. Remarkably, this improvement is achieved despite using synthetic profiles from Germany rather than the UK for pretraining. Fig. 9(b) illustrates that as training duration increases, the benefits of transfer learning diminish—and may even become slightly negative. This is because the model eventually learns all relevant temporal patterns from the target data itself, leaving little advantage to be gained from the pretrained knowledge.
@@
The Transformer model’s accuracy is very stable across varying model sizes, i.e., across different hyperparameter settings. For model sizes of 0.2k parameters or more, the average nMAE fluctuates at most by 0.45%pt, while the standard deviation remains nearly independent of the model’s parameter count, with a maximum fluctuation of 0.27%pt.
@@
All forecasting algorithms are evaluated regarding their forecasting capability for different levels of aggregation. The results are depicted in Fig. 12 for single households (a) and for ECs comprised of 100 households (b).
Figure 12: Load forecasting over (a) 20 single households and (b) 20 communities consisting of 100 households. All other settings follow the baseline configuration.
Fig. 12 demonstrates that the persistence model consistently performs worse than deep learning models at both low and high aggregation sizes. The performance gap is especially noticeable at higher aggregation levels, such as 100 households, where forecast deviations are relatively small.
@@ 3.3 Financial Impact of Forecast Errors
Next to the perfect prediction, which assumes perfect knowledge of the load, the best results are achieved using the Transformer leading to cost savings of up to 8.39%. The remaining two deep learning methods—LSTM and xLSTM—perform comparably well, and the performance differences between the three deep learning methods are minor. It is seen that the difference in average savings between the perfect prediction scenario and the various forecasting scenarios is decreasing with an increased level of aggregation. This is attributed to the improved forecast accuracy for larger, aggregated loads. While deep learning models outperform benchmark models especially in smaller ECs, the performance gap narrows as the community size increases. For large ECs (e.g., 100 households), the relative cost savings converge across all forecasting models, ranging from 8.28% to 8.39%. However, deep learning models still provide a slightly superior load forecast accuracy, as indicated by lower nMAE values.
@@ 4 Conclusions
Although data-driven forecasting methods are gaining widespread attention, their applicability to load forecasting in energy communities has been insufficiently explored. This study addresses that gap through a comprehensive evaluation of state-of-the-art forecasting approaches tailored to energy communities.
Deep learning and conventional models were trained on synthetic communities constructed by randomly sampling load profiles from the UK Power Networks smart meter dataset. A sensitivity analysis was conducted to assess model performance across varying community sizes, training data availability, and model complexity.
While Transformers are often seen as the state-of-the-art for short-term load forecasting, LSTM regression remains a strong competitor in this context. Notably, the newly introduced xLSTM achieves performance comparable to Transformers, offering a robust and viable alternative. Overall, no deep learning architecture demonstrated consistent superiority across all aggregation levels. The model choice can therefore be based on specific needs or preferences, disregarding concerns about significant variations in forecast accuracy.
To support the training of models with limited data, publicly available synthetic load profiles were utilized for transfer learning. This approach demonstrated a significant reduction in forecasting nMAE, particularly in low-data scenarios (e.g., with only two months of data). It underscores the versatility and generalization ability of synthetic profiles for pretraining, especially in cases with limited training data. However, the persistence prediction model was consistently found to outperform all deep learning models when the training data was limited to six months or less—regardless of any prior use of transfer learning.
Beyond predictive accuracy, this study quantifies the financial implications of forecast quality using a mixed-integer linear programming case study. Results show that improved forecast accuracy translates to tangible cost savings for energy communities operating shared battery energy storage systems. For a community of 10 households with a storage capacity of $120~{}\mathrm{kWh}$, savings ranged from 6.7% using a persistence predictor, to 7.2% with xLSTM, and up to 9.0% under perfect forecasting, compared to a baseline without predictive control. Even traditional methods like KNN and persistence forecasting yield significant benefits, despite their lower model complexity and forecast accuracy. This finding is especially relevant for small edge computing devices with limited computational resources, as it suggests that smaller models can achieve comparable performance without the need for extensive computational power.
Future research should expand the scope by comparing forecasting models across a broader range of optimization scenarios. Nevertheless, this work provides valuable insights for operating energy communities and advances the understanding of deep learning methods in this emerging application domain.
@@
Lukas Moosbrugger: Writing – Original Draft, Visualization, Software, Validation, Methodology, Data curation, Formal analysis, Conceptualization. Valentin Seiler: Writing – Review & Editing, Visualization, Validation, Methodology, Conceptualization. Philipp Wohlgenannt: Writing – Review & Editing, Software, Validation, Methodology, Conceptualization. Sebastian Hegenbart: Writing - Review & Editing, Validation, Methodology, Conceptualization. Sashko Ristov: Writing - Review & Editing, Validation, Methodology, Conceptualization. Elias Eder: Writing - Review & Editing, Visualization, Validation, Methodology, Conceptualization. Peter Kepplinger: Writing – Review & Editing, Validation, Conceptualization, Supervision, Resources, Project administration, Funding acquisition.
@@ Appendix C Model Configuration
0.2k Blocks = 1 Heads = 1 Features = 1 sLSTM at 0

### v5

@@ Abstract
Energy communities (ECs) play a key role in enabling local demand shifting and enhancing self-sufficiency, as energy systems transition toward decentralized structures with high shares of renewable generation. To optimally operate them, accurate short-term load forecasting is essential, particularly for implementing demand-side management strategies. With the recent rise of deep learning methods, data-driven forecasting has gained significant attention, however, it remains insufficiently explored in many practical contexts. Therefore, this study evaluates the effectiveness of state-of-the-art deep learning models—including LSTM, xLSTM, and Transformer architectures—compared to traditional benchmarks such as K-Nearest Neighbors (KNN) and persistence forecasting, across varying community size, historical data availability, and model complexity. Additionally, we assess the benefits of transfer learning using publicly available synthetic load profiles. On average, transfer learning improves the normalized mean absolute error by 1.97 percentage points when only two months of training data are available. Interestingly, for less than six months of training data, simple persistence models outperform deep learning architectures in forecast accuracy. The practical value of improved forecasting is demonstrated using a mixed-integer linear programming optimization for ECs with a shared battery energy storage system. For an energy community with 50 households, the most accurate deep learning model achieves an average reduction in financial energy costs of 8.06%. Notably, a simple KNN approach achieves average savings of 8.01%, making it a competitive and robust alternative. All implementations are publicly available to facilitate reproducibility. These findings offer actionable insights for ECs, and they highlight when the additional complexity of deep learning is warranted by performance gains.
keywords:
Load Forecasting , Energy Communities , Deep Learning , Transformer , xLSTM
journal: Next Energy
[…]
@@ 1 Introduction
Energy communities (ECs) are continuously gaining attention, as they accelerate the transition to renewable energy and enhance self-sufficiency by enabling local balancing of supply and demand. ECs are groups of individuals or organizations that produce, manage, and consume energy locally, through the use of public grid infrastructure, often with the goal of increasing sustainability and reducing energy costs [1, 2]. For an efficient operation of these entities—and to improve both self-sufficiency and financial benefit [3]—model predictive control (MPC) methods have been employed with notable success [4, 5].
@@
In 2024, Beck et al. [21] introduced the xLSTM model. The authors claimed that this model performs favorably when compared to state-of-the-art Transformers, both in performance and scaling. They demonstrated that xLSTM exhibits performance comparable to Transformers, with notable advantages such as linear rather than quadratic growth in memory consumption as the model input size increases. This improvement addresses scalability challenges, making xLSTM a more efficient option to handle larger inputs. A first work by Kraus et al. [22] demonstrates the effectiveness of xLSTM for long-term forecasting across various publicly available time series, with forecast horizons ranging from 96 to 720 hours. In contrast to prior work, the present study applies xLSTM in a short-term forecasting context, using a 24-hour horizon aligned with the requirements of our MPC application.
In the context of ECs, the forecast accuracy is often deteriorated by a lack of sufficient historical records [23], irrespective of the used ML-based model. To address this challenge, transfer learning emerges as a promising solution by enabling knowledge reuse from related data. This approach is based on the idea that knowledge learned during one task can be reused to boost the performance on a related task [24]. In Moosbrugger et al. [25], the potential of leveraging transfer learning was demonstrated using open-access synthetic (also called standard) load profiles to improve load forecast error for renewable ECs with limited historical data. The study utilized a bidirectional LSTM model pre-trained on open-access synthetic profiles, achieving a 62% reduction in mean squared error (MSE) for forecasted load in an energy community of 74 households, while also significantly improving training stability.
Even though MPC has been extensively applied in the context of ECs, most studies assume either perfect load forecasts or rely on narrowly defined assumptions regarding the forecasting method. These simplifications limit the applicability of the results, as forecasting models are typically tested in isolated scenarios rather than evaluated across a range of conditions such as community size, training data availability, or model complexity, leaving a gap in understanding how different forecasting algorithms perform under varying conditions. This gap is addressed via the following research questions:
@@
    Is the use of forecasting models based on deep learning methods justified across different EC configurations?
@@
    Can transfer learning using publicly available synthetic load profiles improve the forecast accuracy of deep learning models in data-scarce ECs?
@@
    How is the financial benefit of MPC-based flexibility optimization impacted by the load forecast?
To answer these research questions, load profiles, aggregation levels, training data durations, and model sizes are realistically varied, providing a more comprehensive evaluation. This approach aligns with the recommendation of Pinheiro et al. [26], who noted that many existing studies demonstrate the superiority of a proposed technique on narrowly defined datasets, often hiding their weaknesses. Therefore, this study provides several contributions:
  1. 1.
    We quantify the data availability threshold at which advanced deep learning models begin to significantly outperform conventional baseline models.
  2. 2.
    We propose and evaluate a novel application of transfer learning from publicly available synthetic load profiles for EC load forecasting.
  3. 3.
    We demonstrate the financial implications of forecast accuracy through a case study of a simulated EC, highlighting the economic benefits of using simple models such as k-nearest neighbors (KNN).
This study therefore critically assesses the suitability of current deep learning approaches for household and EC load forecasting by systematically probing their sensitivity to real-world scenarios and investigating the value of improved forecasting accuracy for energy system optimization.
@@ 2 Methods
In the next step, three deep learning models (xLSTM, LSTM, and Transformer) were compared against two baseline models: Persistence Prediction and k-Nearest Neighbors (KNN). Each model performs day-ahead short-term load forecasting, predicting an hourly power load profile from midnight to the following midnight. The models were trained and tested on various configurations in terms of community size, training dataset size, and model complexity. The practical value of improved forecasting was analyzed through a mixed-integer linear programming (MILP) optimization model for ECs with shared battery energy storage.
@@ 2.1 Data description
This study is based on 4,500 household load profiles from the UK Power Networks smart meter dataset [29], which contains electricity consumption data from households in London. To construct virtual ECs, subsets of 1, 2, 10, 50, or 100 households were randomly selected from the dataset. This selection process was repeated 20 times for each community size, without replacement, resulting in $100$ unique virtual ECs. All profiles span the period from July 2012 to February 2014. Q4 of 2013 was used as the test period, with the preceding 12 months used for training in all baseline model evaluations. The effect of using different quarters of 2013 as the testing period was also evaluated in a robustness analysis, which can be found in B.
@@ 2.2.1 Input Features
  • 1.
    Day of the week: Encoded as a one-hot vector, represented as a binary 7-dimensional vector with one ”hot” bit and public holidays assigned to Sunday [15, 32].
  • 2.
@@ 3.2 Sensitivity Analysis
It is seen that the nMAE of the persistent predictor is smaller compared to the deep learning models with up to six months of training data. However, with nine months of training data or more, the deep learning models outperform persistent prediction, with a difference in nMAE ranging from 1.8 percentage points to 2.0 percentage points, depending on the model type. This finding is particularly valuable for newly established ECs, which can initially rely on persistence prediction for MPC and transition to more sophisticated models after operating for nine months.
@@
As illustrated in Fig. 9(a), transfer learning using synthetic load profiles shows a notable reduction in nMAE for two months of available training data. Across all trained models, the average nMAE improvement is 1.97 percentage points. Remarkably, this improvement is achieved despite using synthetic profiles from Germany rather than the UK for pretraining. Fig. 9(b) illustrates that as training duration increases, the benefits of transfer learning diminish—and may even become slightly negative. This is because the model eventually learns all relevant temporal patterns from the target data itself, leaving little advantage to be gained from the pretrained knowledge.
@@
The Transformer model’s accuracy is very stable across varying model sizes, i.e., across different hyperparameter settings. For model sizes of 0.2k parameters or more, the average nMAE fluctuates at most by 0.45 percentage points, while the standard deviation remains nearly independent of the model’s parameter count, with a maximum fluctuation of 0.27 percentage points.
@@
All forecasting algorithms are evaluated regarding their forecasting capability for different levels of aggregation. The results are depicted in Fig. 12 for single households (light blue circles) and for ECs comprised of 100 households (black markers).
Figure 12: Forecast nMAE of 20 single households (light blue circles, $h=1$) and 20 communities consisting of 100 households (orange squares, $h=100$). Circles and squares represent the mean values, while the whiskers indicate the standard deviations. All other settings follow the baseline configuration.
Fig. 12 demonstrates that the persistence model consistently performs worse than deep learning models at both low and high aggregation sizes. The performance gap gets stronger at higher aggregation levels, such as 100 households, where the forecast deviation is small.
@@ 3.3 Financial Impact of Forecast Errors
Next to the perfect prediction, which assumes perfect knowledge of the load, the best results for a 100-household energy community are achieved using the Transformer leading to cost savings of up to 8.39%. The remaining two deep learning methods—LSTM and xLSTM—perform comparably well, and the performance differences between the three deep learning methods are minor. It is seen that the difference in average savings between the perfect prediction scenario and the various forecasting scenarios is decreasing with an increased level of aggregation. This is attributed to the improved forecast accuracy for larger, aggregated loads. While deep learning models outperform benchmark models especially in smaller ECs, the performance gap narrows as the community size increases. For large ECs (e.g., 100 households), the relative cost savings converge across all forecasting models, ranging from 8.28% to 8.39%. However, deep learning models still provide a slightly superior load forecast accuracy, as indicated by lower nMAE values.
@@ 4 Conclusions
This study critically evaluates the suitability of deep learning models for short-term load forecasting in ECs, comparing them against traditional benchmarks such as persistence prediction and KNN regression. The findings suggest that while deep learning methods are often considered state-of-the-art, their performance is not universally superior—especially in scenarios with limited training data.
To improve the performance of the forecasting methods in data-scarce environments, transfer learning from publicly available synthetic load profiles was applied. In scenarios with limited training data—such as only two months—this method reduced forecasting errors by nearly 2 percentage points. This underscores the versatility and generalization ability of these synthetic profiles for pretraining.
Notably, simple methods like persistence forecasting outperform deep learning models when training data is restricted to six months or less, regardless of the use of transfer learning. However, as the training data size increases, deep learning models show significant improvements in forecast accuracy and are getting less reliant on transfer learning.
In addition to predictive performance, the study quantifies the financial implications of load forecasting accuracy using a mixed-integer linear programming case study. For a community of 10 households with a shared battery energy storage system of $120~{}\mathrm{kWh}$, the best-performing deep learning model (xLSTM) achieved average energy cost savings of 7.18%. Interestingly, the simple KNN approach yielded approximately the same cost savings, namely 7.21%, despite its lower model complexity and forecast accuracy.
Thus, the research questions posed in the introduction of this study can be answered as follows:
  1. 1.
    In the context of ECs, deep learning models are not always worth the effort—especially for ECs with less than six months of available data, where simpler models like persistence prediction yield lower forecasting errors.
  2. 2.
    For applications that nevertheless want to use deep learning methods with only a few month of training data, transfer learning from publicly available synthetic load profiles improves the average forecasting error.
  3. 3.
    From a financial perspective, simple methods such as KNN can yield cost savings comparable to deep learning models, making them a viable option especially in resource-constrained environments.
In future research, the results provided by this study can be built on by comparing forecasting models across a broader range of optimization scenarios. Nonetheless, this study contributes valuable insights to the design and operation of energy communities and highlights the trade-offs between model complexity, data availability, and practical benefits. These contributions lay a strong foundation for advancing more efficient and practical energy management strategies.
@@
Lukas Moosbrugger: Writing - Original Draft, Visualization, Software, Validation, Methodology, Data curation, Formal analysis, Conceptualization. Valentin Seiler: Writing - Review & Editing, Visualization, Validation, Methodology, Conceptualization. Philipp Wohlgenannt: Writing - Review & Editing, Software, Validation, Methodology, Conceptualization. Sebastian Hegenbart: Writing - Review & Editing, Validation, Methodology, Conceptualization. Sashko Ristov: Writing - Review & Editing, Validation, Methodology, Conceptualization. Elias Eder: Writing - Review & Editing, Visualization, Validation, Methodology, Conceptualization. Peter Kepplinger: Writing - Review & Editing, Validation, Conceptualization, Supervision, Resources, Project administration, Funding acquisition.
@@ Appendix C Model Configuration
0.2k
Blocks = 1
Heads = 1
Features = 1
sLSTM at 0